# Compositional Hardness of Code in Large Language Models - A Probabilistic Perspective

### v1

@@ 2.3 Effect of noise on decoding
While the noise onto the logits, $X$, averages to zero, its effect on the decoding process does not. The probability of each token in a decoding step is changed to:
$\mathbb{P}(i|context)\rightarrow\mathbb{P}^{\prime}(i|context)\leq\frac{\mathbb{P}(i|context)}{\mathbb{P}(i|context)+(1-\mathbb{P}(i|context))e^{X}}$ (5)
The denominator, $\mathbb{P}(i|context)+(1-\mathbb{P}(i|context))e^{X}$ can be thought of as a renormalizing term, which redistributes the probability of the tokens. For $X=0$, the token’s probability does not change, for $X<0$ it increases and for $X>0$ it decreases. Note that if $\mathbb{P}(i|context)\in(\epsilon,1-\epsilon)$ for $\epsilon>0$ (the model has finite confidence), then on average, the noise decreases the probability $\mathbb{P}(i|context)$, by a factor of $exp(-\Delta(\epsilon,X))$, where $\Delta$ is the renormalizing term’s mean:
@@ 3 Results
Here we show that composing two coding problems can be significantly harder than solving each on its own. A natural quantification of compositional hardness is the gap between generation complexity of the problem’s components and the complete problem. For example, we say that composition is hard if:
$N(\mathbb{P},x)\gg N(\mathbb{P},x_{1})\cdot N(\mathbb{P},x_{2})$ (9)
While it is easy if:
$N(\mathbb{P},x)\approx N(\mathbb{P},x_{1})\cdot N(\mathbb{P},x_{2})$ (10)
The rational behind this, is that $N(\mathbb{P},x_{1})\cdot N(\mathbb{P},x_{2})$ is the number of attempts required to independently sample a correct solution to $x_{1}$ and to $x_{2}$, while $N(\mathbb{P},x)$ is the number of attempts required to sample a solution to problem that integrates both problems. In an easy composition scenario, the model solves the sub-problems to the best of its abilities as it would if each was solved independently. In a hard composition scenario, seeing both problems combined reduces its performance on each sub-problem, and it is better to use a multi-agent system, by feeding the model the subproblems in different contexts and sample solutions independently.
As there are typically more incorrect solutions to coding problems than correct ones, the random noise inserted into the logits generally harms the model’s performance. The following lemma quantifies an exponential decrease in the model’s probability of a correct solution to a compositional problem, $\mathbb{P}(y_{1}\oplus y_{2}|x)$, relative to the probabilties of the sub-problem solutions $\mathbb{P}(y_{1}|x_{1})\cdot\mathbb{P}(y_{2}|x_{2})$:
@@
Let $\epsilon,\delta\in(0,1)$, and $M>0$. Let $x$ be a compositional problem and $y_{1}\oplus y_{2}$ a solution, with $x_{1}$, $x_{2}$ being the corresponding sub-problems. Suppose that the noise injected to the logits as defined in equation 4, satisfies assumption 1 - for all decoding steps, it is continuous, symmetric and bounded within $[-M,+M]$. Suppose further that the probability assigned to the correct token at each decoding step is bounded within $[\epsilon,1-\epsilon]$. Denote by $\Delta:=\mathbb{E}_{X}[\log(\epsilon+(1-\epsilon)e^{X})]$ and $\sigma^{2}:=Var_{X}[\log(\epsilon+(1-\epsilon)e^{X})]$ the renormalizing term’s mean and variance (as defined in equations 6 and 8 respectively). Under the assumption, $\Delta,\sigma$ are strictly positive, and if $|y_{1}|+|y_{2}|>\frac{M^{2}}{\sigma^{2}\cdot h(\frac{3\Delta\cdot M}{4\sigma^{2}})}\ln\frac{1}{\delta}$, where $h(x)=(x+1)\ln(1+x)-x>0$, we have with probability of at least $1-\delta$ that:
$\mathbb{P}(y_{1}\oplus y_{2}|x)\leq\mathbb{P}(y_{1}|x_{1})\cdot\mathbb{P}(y_{2}|x_{2})e^{-\frac{\Delta\cdot(|y_{1}|+|y_{2}|)}{4}}$ (11)
Where $\mathbb{P}(y|x)$ is the probability assigned to $y$ by the model, given context $x$.
@@
The assumption on bounded probability for the correct token $[\epsilon,1-\epsilon]$ implies we are considering solutions where the model has high but limited confidence in each decoding step. While LLMs are often very confident, in practice, when sampling, nucleus sampling is commonly used, where sampling only occurs if the model is not overly confident. e.g. with $p=0.95$, if a token’s probability $P$, is larger than $0.95$ or much smaller than $0.05$, then the probability is rounded to 00 or $1$. Thus when considering probabilities of sequences, it suffices to look only at decoding steps where the model is not too confident, hence we can consider $\epsilon\approx 0.05$.
In subsection 4.3, we see empirically that for $\epsilon=0.1$, $\Delta\approx 0.2,\sigma\approx 1.5,M\approx 4$, making $\frac{M^{2}}{\sigma^{2}\cdot h(\frac{3\Delta\cdot M}{4\sigma^{2}})}\approx 200$, thus for solutions with length $>200$ tokens, the results apply.
In practice, there may be multiple solutions to the same problem, e.g. multiple implementations of the same function. So it is necessary to take all of them into account when considering the generation complexity. With this taken into account using a union bound, we obtain the following result of an exponential gap in generation complexity between a composition of problems and the sub-problems, indicating a compositional hardness that is exponential in the solution’s length:
@@
Let $\epsilon,\delta\in(0,1)$, and $N,M>0$. Let $x$ be a compositional problem, with $x_{1}$, $x_{2}$ being the corresponding sub-problems. Denote by $L_{1},L_{2}$ the minimal solution length to $x_{1},x_{2}$ respectively, and the total number of solutions to $x$ by $N$. Define $\Delta,\sigma$, the renormalizing term’s mean and variance (as defined in equations 6 and 8 respectively) and by $M$ the bound on the logit noise (assumption 1). Under the assumptions of lemma 1, they are strictly positive, $\Delta,\sigma,M>0$, and if the minimal solution length $L_{1}+L_{2}$, satisfies $L_{1}+L_{2}>\frac{M^{2}}{\sigma^{2}\cdot h(\frac{3\Delta\cdot M}{4\sigma^{2}})}\ln\frac{N}{\delta}$, where $h(x)=(x+1)\ln(1+x)-x>0$, we have with probability of at least $1-\delta$ that the generation complexity (definition 2) satisfies:
$N(\mathbb{P},x)\geq N(\mathbb{P},x_{1})N(\mathbb{P},x_{2})\cdot e^{\frac{\Delta\cdot(L_{1}+L_{2})}{4}}$ (12)
The proof is presented in appendix B. We see that longer problems become harder to solve due to the noise injected into the decoding steps by previously generated tokens. This shows that a model that fully utilizes its context in decoding (i.e. the next token probability distribution is explicitly a function of all the previous tokens), can have a hard time mixing different concepts due to the screening effect. This result implies that for long coding problems, it is more beneficial to distribute sub-tasks between different instances of the LLM, and not expose it to the full context. Additionally, this result provides a view of the model’s effective context length through the lens of screening – additional irrelevant context may reduce model’s performance on other tasks within the context window exponentially with length. The intuition to the result is that there are more incorrect solutions than correct ones, so the random noise usually reduces their total probability mass.
@@
We note that for the union bound of theorem 1 to hold, we need the number of solutions to be bounded by an exponential in the length of the solution, $L_{1}+L_{2}$:
$N<\delta\cdot exp\left((L_{1}+L_{2})\cdot\frac{\sigma^{2}}{M^{2}}h\left(\frac{3\Delta\cdot M}{4\sigma^{2}}\right)\right)$ (13)
@@
Typically, we expect the number of solutions to a problem to grow exponentially with the length of the shortest solution, $y$:
$N\sim exp(c\cdot|y|)$ (14)
@@
As each line of code can be implemented slightly differently, lines may be interchanged, different variable names, etc. Still, the exponential’s coefficient is expected to be small as most sequences are not correct solutions to the problem due to the constraints between the tokens (e.g., once a variable name is chosen, it is fixed throughout the solution). With the empirical values calculated in subsection 4.3, we see empirically that the coefficient of the exponential number of solutions is $\approx 0.005$, thus for a solution of length $L_{1}+L_{2}=1,600$ tokens, we have $N<\delta\cdot exp(8)=3,000\cdot\delta$ solutions.
@@ 4 Experiments
In this subsection we test the assumptions and results of our theoretical part. We create simple compositional coding problems with pairs of problems from the humaneval benchmark Chen et al. (2021a) and code contests dataset Li et al. (2022b). First, we show the results of theorem 1, stating that composition is typically harder than independently solving the subproblems. Then, we show explicit indications for the exponential length dependence of compositional hardness, by comparing probabilities of solutions with/without composition, as theoretically suggested in lemma 1. Finally, we look at our assumption 1 on the noise inserted into the logits of the second problem, and observe that it is indeed large enough to interfere with the decoding process (assumption 1). The experiments were performed on Llama-3-8B-Instruct Dubey et al. (2024).
@@ 4.1 Generation Complexity Results
Here we test the actual generation complexity of an LLM to different problems corresponding to scenarios from our theoretical results. As proposed in the theoretical section, an LLM may be able to solve two problems with low generation complexity, but their composition might take a significantly larger generation complexity if the model was not explicitly trained on such a task. To test this, we built a set of composite problems based on the Human Eval benchmark Chen et al. (2021a) and code contests dataset Li et al. (2022b).
@@
To create composite problems, we took pairs of problems from Human Eval and code contests, and created from each a problem whose solution requires to explicitly solve both problems. Additionally, for harder problems, we created compositions of problems from code contests dataset. The three templates used were:
  • (Human Eval) Problem 1 has a True/False output, Problem 2 has an arbitrary output. Composition is to solve problem 1, then if its output is true, print the second problems’ output, otherwise, print "-1":
@@
  • (Human Eval) Problem 1 and Problem 2 have an integer output. Composition is to solve both and print the product of their outputs.
@@
  • (Code Contests) Problem 1 and problem 2 are problems from the dataset. Composition is to read the inputs sequentially and print the outputs sequentially.
We denote the composite problem of $x_{1}$ and $x_{2}$ as $x_{1}\oplus x_{2}$. For each problem we sampled 200 solutions and evaluated the generation complexity as the inverse of percentage of correct solutions.
@@
Figure 1 shows the cumulative distribution function (CDF) of the compositional hardness - the value along the $y$ axis is the CDF of $\frac{N(\mathbb{P},x)}{N(\mathbb{P},x_{1})N(\mathbb{P},x_{2})}$, i.e.  percentage of compositional problems, in which $\frac{N(\mathbb{P},x)}{N(\mathbb{P},x_{1})N(\mathbb{P},x_{2})}$ is smaller than a given value. For example, with $a=5$, the corresponding value on the $y$ axis, is the percentage of problems in which composition requires up to ×$\times 5$ more generations relative to solving independently. As can be seen, for most of the problems, composition generation complexity is larger than the product of generation complexities of the components (up to factors of $10-20$).
@@
Figure 1: Cumulative distribution function for the ratio of generation complexity using composition, $N(\mathbb{P},x)$, to product of generation complexities for the standalone problems, $N(\mathbb{P},x_{1})\cdot N(\mathbb{P},x_{2})$ (corresponding to the multi-agent generation complexity). The x axis denotes values for the ratio of generation numbers required to solve the problem in the two cases (composition vs multi-agent), the y axis is the percentage of problems in which the ratio is no larger than this value (e.g. for $a=5$, the y axis value is the percentage of problems where less than ×$\times 5$ samples are required in composition than multi-agent). (a) For the human eval composition. As can be seen in most of the cases, composition requires twice more samples, and for some problems 10 times more samples. (b) For the code contests composition. As can be seen the majority of problems have a factor of at least 5, and some up to 20.
@@
$N(\mathbb{P},x_{1}\oplus x_{2})\gg N(\mathbb{P},x_{1})N(\mathbb{P},x_{2})$ (15)
Furthermore, we see that in code contests, where the problems typically have longer solutions (few hundreds of tokens), and the problems are harder thus the model is less certain, the compositional hardness is much larger.
Assuming one only has access to end-to-end verifier of the composite problem, it is more advantageous to generate independently for each problem then to generate for the composite problem, giving the advantage to a multi-agent system over a single LLM instance.

### v3

@@ 3.3 Effect of noise on decoding
While the noise onto the logits, $X$, averages to zero, its effect on the decoding process does not. The probability of each token in a decoding step given context $c$ is changed to:
$P(i|c)\rightarrow P^{\prime}(i|c)\leq\frac{P(i|c)}{P(i|c)+(1-P(i|c))e^{X}}$ (8)
The denominator, $P(i|c)+(1-P(i|c))e^{X}$ can be thought of as a renormalizing term, which redistributes the probability of the tokens. For $X=0$, the token’s probability does not change, for $X<0$ it increases and for $X>0$ it decreases. See appendix E for derivation and intuition. Note that if $P(i|c)\in(\epsilon,1-\epsilon)$ for $\epsilon>0$ (the model has finite confidence), on average, the noise decreases the probability of a correct continuation $P(i|c)$, by a factor $\exp(-\Delta(\epsilon,X))$, where $\Delta$ is the renormalizing term’s mean:
@@ 4 Results
Here we show that composing coding problems can be significantly harder than solving them in parallel contexts. This helps us derive generation complexity bounds for general problems in terms of simpler problems the model knows how to solve. We present results for composition of two problems, but note it can inductively be increased to any number. A natural quantification of compositional hardness is the ratio between generation complexity of the problem’s components and the complete problem.
$\frac{N(P,(x_{1},x_{2}))}{N(P,x_{1})\cdot N(P,x_{2})}$ (12)
A composition is hard if this ratio is significantly larger than $1$, while easy if it is close to $1$. The rational, is that $N(P,x_{1})\cdot N(P,x_{2})$ is the number of attempts required to sample in parallel a correct solution to $x_{1}$ and $x_{2}$, while $N(P,(x_{1},x_{2}))$ is the number of attempts required to sample a solution to both in a single context. When composition is easy, the model solves the each sub-problem to the best of its abilities, but when composition is hard, seeing both problems combined reduces its performance on each sub-problem, and a multi-agent approach is more favorable.
As there are typically more incorrect solutions to coding problems than correct ones, the random noise inserted into the logits generally harms the model’s performance. The following lemma quantifies an exponential decrease in the model’s probability of a correct solution to a compositional problem, $P(y_{1}\oplus y_{2}|x)$, relative to the probabilties of the sub-problem solutions $P(y_{1}|x_{1})\cdot P(y_{2}|x_{2})$:
@@
Let $\epsilon,\delta\in(0,1)$, and $M>0$. Let $x$ be a compositional problem and $y_{1}\oplus y_{2}$ a solution, with $x_{1}$, $x_{2}$ being the corresponding sub-problems. Suppose that the noise injected to the logits as defined in equation 7, satisfies assumption 1, and that the probability assigned to the correct token at each decoding step is bounded within $[\epsilon,1-\epsilon]$. Then there exist strictly positive noise dependent constants $\Delta$ (as defined in equation 9) and $c(\Delta,M,\sigma)$ (with $M$ and $\sigma$ as defined in equations 10 and 11), such that if the solution length satisfies $|y_{1}|+|y_{2}|>c\ln\frac{1}{\delta}$ we have with probability of at least $1-\delta$ (over noise randomness) that:
$P(y_{1}\oplus y_{2}|x)\leq P(y_{1}|x_{1})\cdot P(y_{2}|x_{2})e^{-\frac{\Delta\cdot(|y_{1}|+|y_{2}|)}{4}}$ (13)
Where $P(y_{1}\oplus y_{2}|x)$ is the probability of producing the answer $y_{1}\oplus y_{2}$, given context $x$. The constant $c=\frac{M^{2}}{\sigma^{2}\cdot h(\frac{3\Delta\cdot M}{4\sigma^{2}})}$, with $h(x)=(x+1)\log(x+1)-x$.
@@
The assumption on bounded probability for the correct token $[\epsilon,1-\epsilon]$ implies we are considering solutions where the model has high but limited confidence in each decoding step. While LLMs are often very confident in “obvious" next steps (line break, etc.), in practice, during generation, nucleus sampling is commonly used, where sampling only occurs if the model is not overly confident. e.g. with $p=0.95$, if a token’s probability $P$, is larger than $0.95$, then the probability is rounded $1$. Thus when considering probabilities of sequences, it suffices to look only at decoding steps where the model is not too confident, hence we can consider $\epsilon\approx 0.05$.
In subsection 5.3, we see empirically that for $\epsilon=0.1$, $\Delta\approx 0.2,\sigma\approx 1.5,M\approx 4$, making $c\approx 200$, thus for solutions with length $>200$ tokens, the results apply.
In practice, there may be multiple solutions to the same problem (e.g. different implementations). So it is necessary to take all of them into account when considering the generation complexity. With this taken into account using a union bound, we obtain the following result of an exponential gap in generation complexity between a composition of problems and the sub-problems, indicating a compositional hardness that is exponential in the solution’s length:
@@
Let $\epsilon,\delta\in(0,1)$, and $N,M>0$. Let $x$ be a compositional problem, with $x_{1}$, $x_{2}$ being the corresponding sub-problems. Denote by $L_{1},L_{2}$ the minimal solution length to $x_{1},x_{2}$ respectively, and the total number of solutions to $x$ by $N$. Under the assumptions of lemma 1, there exist strictly positive noise dependent constants $\Delta$ (as defined in equation 9) and $c(\Delta,M,\sigma)$ (with $M$ and $\sigma$ as defined in equations 10 and 11), such that if the minimal solution length $L_{1}+L_{2}$, satisfies $L_{1}+L_{2}>c\ln\frac{N}{\delta}$, then with probability of at least $1-\delta$ over the noise’s randomness, the generation complexity (definition 2) satisfies:
$N(P,x)\geq N(P,x_{1})N(P,x_{2})\cdot e^{\frac{\Delta\cdot(L_{1}+L_{2})}{4}}$ (14)
The proof is presented in appendix B. We see that longer problems become harder to solve within the same context relative to solving them in parallel due to the noise injected into the decoding steps by previously generated tokens. We reiterate this is a stronger statement than the success rate dropping exponentially with the number of compositions, as shown in Dziri et al. (2024), but rather that success rate drops exponentially relative to solving the problems in parallel (which in itself is also exponential in number of compositions), leading to a multi-agent efficiency. This result implies that for long coding problems, it is more beneficial to distribute sub-tasks between different instances of the LLM, and not expose it to the full context.
@@
We note that for the union bound of theorem 1 to hold, the number of solutions needs to be bounded by an exponential in the length of the solution, $L_{1}+L_{2}$:
$N<\delta\cdot\exp\left((L_{1}+L_{2})\cdot\frac{\sigma^{2}}{M^{2}}h\left(\frac{3\Delta\cdot M}{4\sigma^{2}}\right)\right)$ (15)
@@
Typically, we expect the number of solutions to a problem to grow exponentially relative to shortest solution length, $y$:
$N\sim\exp(c\cdot|y|)$ (16)
@@
As variable names may be changed, lines implemented differently, etc. Still, the exponential’s coefficient is small as most sequences are not correct solutions due to constraints between the tokens (e.g., once a variable name is chosen, it is fixed throughout the solution). With the empirical values calculated in subsection 5.3, we see empirically that the coefficient of the exponential number of solutions is $\approx 0.005$, thus for a solution of length $L_{1}+L_{2}=1,600$ tokens, we have $N<\delta\cdot\exp(8)=3,000\cdot\delta$ solutions.
4.1 Generation Complexity For General Problems
Based on the above results for compositional problems, we derive the following lower bound on general code problems for which multiple implicit decompositions may exist:
Theorem 2.
Let $x$ be a coding problem. Suppose the model generates a solution by 1) Sampling a decomposition candidate $x_{1}...x_{k}$ using CoT, 2) Sampling a solution to the compositional problem $y_{1}\oplus...\oplus y_{k}$. Denote by $\Delta$ and $c$ the constants from theorem 1 and assume its conditions hold. Denote $d(x)$ the set of all valid decompositions to the problem for which $P(x_{1}...x_{k}|x)>0$ (i.e. those the model can generate). If for all $x_{1}...x_{k}\in d(x)$ the minimal solution length $L_{1}...L_{k}$ satisfies $\sum_{i=1}^{k}L_{i}>c\ln(\frac{N\cdot|d(x)|}{\delta})$, where $N$ is the number of solutions to the decomposition, then with probability $1-\delta$, the generation complexity for $x$ is lower bounded by:
$N(P,x)\geq[\Pi_{i=1}^{k}N(P,x^{*}_{i})]\cdot exp(\Delta\cdot\sum_{i=1}^{k}L_{i})$ (17)
Where $x^{*}_{1}...x^{*}_{k}$ is the lowest generation complexity decomposition in $d(x)$ and $L_{1}...L_{k}$ is its minimal solution length.
This bounds the success rate of solving a general problem in terms of simpler tasks that the model may choose to decompose the problem to. Proof presented in appendix D.
@@ 5 Experiments
In this subsection we test the assumptions and results of our theory. We create simple compositional coding problems with pairs of problems from the Human Eval benchmark (Chen et al., 2021a) and code contests dataset (Li et al., 2022b). The choice to use pairs of problems and not more, is the low pass rate of $\sim 1/200$ of composition, requiring many samples. First, we show the results of theorem 1, stating that composition is typically harder than solving the the subproblems in parallel. Then, we show explicit indications for the exponential length dependence of compositional hardness, by comparing probabilities of solutions with/without composition, as theoretically suggested in lemma 1. Finally, we look at our assumption 1 on the noise inserted into the logits of the second problem, and observe that it is indeed large enough to interfere with the decoding process (assumption 1). The experiments were performed on Llama-3-8B-Instruct (Dubey et al., 2024). In appendix G we show the results apply also to Mistral-7B-Instruct-v0.3 (Jiang et al., 2023), and in appendix H, we present results for Llama-3-70B-Instruct, to test dependence on model size. For additional experimental details, see appendix F.
@@ 5.1 Generation Complexity Results
Here we test the actual generation complexity of an LLM to different problems corresponding to scenarios from our theoretical results. As proposed in the theoretical section, the generation complexity to solve two problems in a single context can be much higher than that of parallel contexts if the model was not explicitly trained on such a task. To test this, we built a set of composite problems based on the Human Eval benchmark (Chen et al., 2021a) and code contests dataset (Li et al., 2022b).
@@
To create composite problems, we took pairs of problems from Human Eval and code contests, and created from each a problem whose solution requires to explicitly solve both problems. Additionally, for harder problems, we created compositions of problems from code contests dataset. We used the following two main templates (and an additional template for human eval presented in appendix F):
@@
  • Human Eval – Problem 1 and Problem 2 have an integer output. Composition is to solve both and print the product of their outputs.
@@
  • Code Contests – Problem 1 and problem 2 are problems from the dataset. Composition is to read the inputs sequentially and print the outputs sequentially.
We denote the composite problem of $x_{1}$ and $x_{2}$ as $x_{1}\oplus x_{2}$. While these compositions are synthetic, their purpose is to demonstrate that the LLM suffers from a “mental load" when asked to solve multiple tasks within the same context, similarly to a problem decomposition obtained from a CoT, which leads to the result of theorem 1. For each problem we sampled 200 solutions and evaluated the generation complexity as the inverse of percentage of correct solutions.
@@
Figure 1 shows the cumulative distribution function (CDF) of the compositional hardness – the value along the $y$ axis is the CDF of $\frac{N(P,x)}{N(P,x_{1})N(P,x_{2})}$, i.e.  percentage of compositional problems, in which $\frac{N(P,x)}{N(P,x_{1})N(P,x_{2})}$ is smaller than a given value. e.g., $x=5$, the value on the $y$ axis, is percentage of problems in which composition requires up to ×$\times 5$ more generations relative to solving independently. As can be seen, for most of the problems, composition generation complexity is larger than the product of generation complexities of the components (up to factors of $10-20$).
@@
Figure 1: Cumulative distribution function for the ratio of generation complexity using composition, $N(P,x)$, to product of generation complexities for the standalone problems, $N(P,x_{1})\cdot N(P,x_{2})$ (corresponding to the multi-agent generation complexity). The x axis denotes values for the ratio of generation complexities, the y axis is the percentage of problems in which the ratio is no larger than this value (e.g. for $a=5$, the y axis value is the percentage of problems where composition requires up to ×$\times 5$ more samples than the multi-agent case). (a) For the human eval composition. As seen, in most cases composition requires twice more samples, and sometimes up to 10 times more samples. (b) For the code contests composition. As seen, the majority of problems have a factor of at least 5, and some up to 20.
@@
$N(P,x_{1}\oplus x_{2})\gg N(P,x_{1})N(P,x_{2})$ (18)
Furthermore, in code contests, where the problems typically have longer solutions (few hundred tokens) and are harder, thus the model is less certain, the compositional hardness is much larger. If one only has access to an end-to-end verifier of the composite problem, it is more advantageous to generate independently for each problem then to generate for the composite problem, giving the advantage to a multi-agent system over a single LLM instance.
Similar results are shown for Mistral-7B-Instruct-v0.3 in appendix G In appendix H, we present results for composition on Llama-3-70B-Instruct, and find that on the same code contests compositions, the compositional generation complexity improves significantly relative to its 8B counterpart. However, with slightly harder compositions of code contest problems involving four problems instead of two, the 70B model’s performance on composition drops significantly, with most compositions requiring over 5-10 times more generations than the non-compositional case. This hints larger models are more efficient at composition yet still suffer from this effect as compositional difficulty rises.